OpenStudy (ashleyisakitty):

The volumes of two similar solids are 1728 and 343. The surface area of the larger solid is 576. What is the surface area of the smaller solid?

OpenStudy (ashleyisakitty):

**AFK will be back in 10

OpenStudy (compassionate):

This will be my last answer before I do my own work.

OpenStudy (ashleyisakitty):

back

OpenStudy (ashleyisakitty):

Thank you Stephen :3 Youre helping a lot

OpenStudy (anonymous):

I think if they're similar solids, the ratio of their volumes is the same as the ratio of their surface areas.

OpenStudy (compassionate):

ITT: @bagajr attempts to snipe me. Give me a moment, Ash, working it out on paper.

OpenStudy (anonymous):

Oh sorry lol

OpenStudy (compassionate):

Scale factor = (1728/343)^(1/3) = 12/7 Let the surface area of the smaller solid be A, then: A * (12/7)^2 = 576 So A = 196 I'm off to do my own work now, Ashley. Good luck on the rest of your exams. Merry Christmas.

OpenStudy (ashleyisakitty):

<3333 thank youuuuuu